### v1

@@ 4.5.6 Model Component Analysis
‘w/o’ represents without.
‘SF’ represents the Social Context Feature Module.
‘ML’ represents the Multimodal Graph Learning Module
‘FI’ represents the Feature Extraction Interaction Module
CrisisSpot (w/o SCFM): The model that excludes the Social Context Feature module, comprises only two modules: FEIM and MGLM. Within this configuration, we attain an accuracy of 94.63%, a precision of 95.86%, a recall of 96.42%, and an F1-score of 96.14%. Here, both textual and visual features interacted via inverted dual embedded attention, which captures complex patterns as discussed in subsubsection 3.1.3. In MGLM, due to graph knowledge propagation, the features extracted from the multimodal encoder get enriched, hence improving the performance of the model.
CrisisSpot (w/o MGLM): The model without the multimodal graph learning module is denoted as the CrisisSpot (w/o MGLM). In this, we achieve 93.98%, 94.20%, 95.98%, and 94.08% as accuracy, precision, recall, and F1-score. Here, the Social Contextual Features enhance the contextual understanding of the model, and due to the interaction of textual and visual modalities in FEIM, the performance of the model gets enhanced.
CrisisSpot (w/o FEIM): The model that excludes the feature extraction and interaction module is labeled as the CrisisSpot (w/o FEIM). In this configuration, we achieve an accuracy of 95.32%, a precision of 95.44%, a recall of 97.44%, and an F1-score of 97.26%. This model is deprived of the attention mechanism, text features, and visual features, due to the exclusion of the FEIM module. In this model, with the help of multimodal integration of features extracted from the multimodal encoder, text-centric scores, user-centric scores, and graph learning, the model significantly outperforms the other variants.
CrisisSpot (FEIM + MGLM + SCFM): This represents our proposed model, which outperforms all other variations, as shown in Fig. 4. This underscores the efficacy of our proposed approach in comparison to the other variants.
@@ 5 Conclusion
In this article, we have introduced an innovative approach for the classification of disaster-related content during emergency situations. Social media platforms have emerged as real-time information sources during disasters; hence, through this article, we highlight both their potential benefits and challenges. By addressing these challenges, our work contributes significantly to disaster response and public safety efforts. Our primary contribution lies in the development of a multimodal attentive learning framework that leverages graph-based and attention-based learning techniques, enhancing the model’s ability to discern complex relationships and subtle nuances within disaster-related content. We have proposed an attention mechanism, Inverted Dual Embedded Attention (IDEA), to capture both harmonious and contrary information within the content. Our proposed method effectively harnesses multimodal data and Social Context Features to create high-level feature representations, significantly improving model performance. This study concludes that the Crisis Informative Score can be valuable in the identification and categorization of disaster-related content. Furthermore, the research demonstrates that User Informative Score and User Engagement metrics can contribute to assessing the credibility of information shared on social media platforms. Due to the dearth of available training corpus in the disaster domain, we have created a novel multimodal crisis dataset, TSEqD (Turkey-Syria Earthquake Dataset), and annotated for both Informative and Humanitarian Tasks. The novel dataset may prove to be a valuable resource for future research in this domain. Through extensive experiments on two datasets, we demonstrate that CrisisSpot surpasses state-of-the-art benchmarks by substantial margins. In essence, the findings of our work may help researchers build a system for emergency responders and humanitarian organizations utilizing disaster-related social media data.

### v2

@@ 5.5.4 Performance Gain Analysis
5.5.4.1 Analyzing Combinations of Modalities
We systematically evaluate the performance of CrisisSpot across a range of modalities combinations. Given the diverse characteristics of each modality, it is crucial to examine the impact of different modality combinations employed in our approach. Our method integrates inputs from three distinct modalities: textual (T), visual (I), and Social Context Features (SCF). The performance assessments of these various modality combinations are elaborated in Table 8.
To evaluate the importance of image input, we exclude it while preserving text and Social Context Features, resulting in CrisisSpot (w/o Images). However, this exclusion may impose limitations on the model. Firstly, the absence of images results in a loss of crucial visual context, impairing the model’s ability to comprehensively understand the data. Additionally, without integrating images, the model misses out on the synergistic benefits derived from combining text and visual data. This leads to a restricted understanding of the data, hindering its performance. Furthermore, the absence of images prevents the utilization of attention mechanisms, limiting the model’s ability to focus on relevant content. The lack of cross-modal feature interaction further restricts the model’s effectiveness. Consequently, CrisisSpot (w/o Images) may face challenges in achieving optimal performance compared to the proposed model that incorporates all modalities.
The CrisisSpot (w/o Text) approach, designed to assess the significance of textual data, excludes text while retaining images and Social Context Features as inputs. By excluding text, we aim to analyze the impact of textual information on the model’s performance. This enables us to discern the specific role of text in enhancing the understanding and classification of crisis-related content. CrisisSpot (w/o Text) presents several technical implications. Firstly, without textual information, the model may struggle to grasp the semantic context embedded within crisis-related content, potentially leading to misinterpretations or incomplete understanding. Additionally, text often provides valuable contextual cues crucial for accurately categorizing crisis-related data. The absence of text input deprives the method of this contextual information, reducing its ability to make informed decisions or classifications. Without text modality, the model may struggle to capture cross-modal relationships effectively, limiting its ability to leverage complementary information from both modalities. Consequently, the exclusion of text may result in decreased classification accuracy and a loss of discriminative power, as textual data often contributes discriminative features essential for accurate classification. In summary, CrisisSpot (w/o Text) may experience reduced semantic understanding, limited contextual information, inability to capture text-image relationships, and decreased classification performance due to the absence of textual data. These technical implications underscore the critical importance of incorporating textual information into multimodal methods for effective crisis content classification.
CrisisSpot (w/o SCF) focuses on evaluating the significance of Social Context Features (SCF) by excluding them while preserving textual and visual inputs. This configuration enables an assessment of the individual contributions of SCF to CrisisSpot’s overall performance. SCF comprises five factors: User Information Source (UIS), Crisis Information Source (CIS), user engagement history, emotional analysis, and sentiment analysis. CrisisSpot (w/o SCF) would lack crucial information regarding the social context surrounding crisis-related content. Specifically, without incorporating factors such as User Information Source (UIS), Crisis Information Source (CIS), user engagement history, emotional analysis, and sentiment analysis, the method would miss out on valuable insights into the credibility, relevance, and sentiment of the content. This omission could result in several consequences. Firstly, the CrisisSpot would have a limited understanding of the social context surrounding crisis events, potentially leading to misinterpretations or incomplete assessments of the severity and urgency of the situation. Additionally, without considering factors such as, user engagement history and sentiment analysis, the method may struggle to accurately gauge the relevance and impact of crisis-related content, affecting its ability to prioritize and respond effectively. Moreover, the absence of social context scores may hinder the model’s ability to make informed decisions regarding the credibility and trustworthiness of information. Furthermore, without user-centric scores and other social context features, CrisisSpot (w/o SCF) may overlook valuable opportunities for engaging with affected communities or leveraging user-generated content to enhance situational awareness and response efforts. In summary, omitting social context scores from CrisisSpot (w/o SCF) would result in a diminished ability to accurately assess the social context surrounding crisis events, potentially compromising the effectiveness of the model in disaster management tasks.
CrisisSpot (T+I+SCF), integrating textual modality, visual modality, and Social Context Features (SCF), demonstrates superior performance compared to all other variants across various evaluation metrics as can seen in Table 8. This enhancement in CrisisSpot (T+I+SCF) compared to its variants stems from the effective utilization of multimodal information, cross-modal relationships, attention mechanisms, and Social Context Features inherent in the posts. These results underscore the efficacy of our proposed multimodal approach and highlight the importance of leveraging each modality integrated into the model for achieving superior performance in crisis content classification tasks.
5.5.4.2 Analysis of Attention Mechanisms
We conduct a comprehensive analysis of different attention mechanisms in CrisisSpot, ranging from methods without attention to those employing classical attention mechanisms such as additive attention, self-attention, and cross-attention. All methods are trained and evaluated on the CrisisMMD dataset using four key metrics: accuracy, precision, recall, and F1-score. We gain crucial insights into the performance of these attention mechanisms. Our findings reveal that compared to the proposed method without attention, all classical attention mechanisms exhibit notable enhancement. This highlights the significance of directing the model’s focus toward relevant information, thereby enhancing CrisisSpot’s ability to identify crisis effectively. However, the most remarkable performance is demonstrated by our proposed IDEA (Inverted Dual Embedding Attention) model. CrisisSpot with IDEA outperforms all compared methods across all metrics, showcasing a remarkable recall of 99.80%. This suggests that IDEA excels in capturing a higher proportion of actual crisis content, due to its utilization of inverted attention to filter out irrelevant information. Moreover, IDEA maintains a high precision of 96.70%, indicating its proficiency in minimizing false positives. This robust performance underscores the effectiveness of IDEA in accurately identifying crisis data. In summary, our ablation study confirms the vital role of attention mechanisms in CrisisSpot for disaster content classification. It underscores the substantial contribution of the proposed IDEA mechanism, which leverages inverted attention and dual embedding spaces to achieve superior performance compared to classical attention mechanisms. These findings provide compelling evidence for the value of IDEA in enhancing CrisisSpot’s performance and its potential to achieve a more accurate and comprehensive understanding of crisis content.
5.5.4.3 Analysis of Social Context Features
In this section, we delve into the intricate role of Social Context Features (SCF) within CrisisSpot, focusing on User Informative Score (UIS), Crisis Informative Score (CIS), and other related features as depicted from Table 8. These features play a pivotal role in understanding the context and informativeness of social media content during the crisis. CrisisSpot (w/o UIS) may struggle to discern the credibility and significance of user-contributed content during a crisis. UIS serves as a crucial indicator of content informativeness, allowing the model to prioritize credible and informative content over noise and misinformation. Without UIS, CrisisSpot’s ability to distinguish between reliable and less informative content may be compromised, potentially leading to an increase in false positives and reduced overall accuracy in crisis detection. Additionally, the absence of a Crisis Informative Score poses a significant challenge in identifying crisis-related vocabulary within textual data. CIS serves as a key indicator of the presence of crisis-relevant terms, enabling CrisisSpot to discern the informativeness of crisis-related content. CrisisSpot (w/o CIS) may struggle to capture essential crisis signals embedded within textual data, leading to potential misclassifications or oversights in crisis classification. As a result, the model’s overall precision, recall, and F1-score may suffer, impacting its effectiveness in crisis content classification. In Table 8, for simplicity, we denote CrisisSpot (w/o other features) as CrisisSpot (w/o OF). Neglecting additional features such as Sentiment Quotient, Emotion Quotient, and User Engagement Metrics deprives CrisisSpot of valuable contextual information necessary for understanding the emotional context and relevance of crisis-related content. Emotion and sentiment analysis are crucial for interpreting the underlying sentiments and nuances expressed in crisis-related posts, while user engagement metrics provide insights into the relevance and impact of content within the crisis context. Without considering these features, CrisisSpot may overlook important interactions, emotional cues, and contextual nuances present in crisis-related content, potentially
leading to suboptimal performance in crisis classification. Our proposed, CrisisSpot (CIS + UIS + Other Features), integrates all SCF components, yielding superior results across all evaluation metrics. By incorporating UIS, CIS, and other related features, CrisisSpot achieves a comprehensive understanding of crisis-related content, enabling more accurate and reliable crisis classification. The inclusion of these features enhances CrisisSpot’s ability to discern credible, informative, and contextually relevant content amidst the noise of social media during crises, underscoring the significance of SCF integration in crisis management and response.
5.5.4.4 Model Component Analysis
In this section, we delve deeply into the intricate components of our CrisisSpot model, dissecting their individual contributions and collective impact on performance, as illustrated in Table 8. Through a series of ablation experiments, we systematically explore the role of each module, shedding light on their significance in crisis content classification. The various variants are as:
Our first variant, CrisisSpot (w/o SCFM), excludes the SCFM, focusing solely on the FEIM and MGLM modules. Despite the absence of SCFM, which typically provides invaluable contextual insights, this configuration remarkably maintains competitive performance metrics. The effective interaction between textual and visual features facilitated by FEIM, coupled with knowledge propagation through MGLM, underscores the model’s ability to capture nuanced patterns within crisis-related content. However, without SCFM, the model may lack deep contextual understanding derived from social context features, potentially leading to suboptimal performance in scenarios where such information is critical.
Next, we evaluate the impact of the MGLM, CrisisSpot (w/o MGLM), by excluding it from the model architecture, while retaining SCFM and FEIM. Despite the omission of MGLM, known for enriching feature representation through graph knowledge propagation, our model exhibits significant performance. The contextual understanding provided by SCFM, coupled with the effective interaction between textual and visual modalities facilitated by FEIM, underscores the model’s resilience in crisis content classification. However, the absence of MGLM may limit the model’s ability to exploit complex relationships between different modalities, potentially resulting in a less nuanced understanding of crisis-related content.
The next variant, CrisisSpot (w/o FEIM), removes the FEIM module from the model, depriving it of attention mechanisms and modality-specific features. Despite this limitation, the model demonstrates commendable performance, highlighting the efficacy of multimodal integration and contextual understanding provided by SCFM and MGLM. Leveraging features extracted from the multimodal encoder, along with text-centric and user-centric scores, the model compensates for the absence of FEIM, showcasing competitive accuracy, precision, recall, and F1-score. However, without FEIM, the model may struggle to capture fine-grained patterns and relationships within and between modalities, potentially leading to reduced performance in scenarios requiring a nuanced understanding of crisis-related content.
The proposed, CrisisSpot (FEIM + MGLM + SCFM), method integrates all components, synergistically leveraging attention mechanisms, modality-specific features, and SCFs to achieve superior performance. By combining FEIM, MGLM, and SCFM, CrisisSpot surpasses other variants, demonstrating the effectiveness of our holistic approach. The seamless integration of these modules enables the model to capture nuanced patterns within crisis-related content, resulting in significant accuracy, precision, recall, and F1-score. Here, the absence of any individual component may significantly diminish the model’s overall performance, underscoring the importance of comprehensive multimodal integration and SCFs in crisis content classification. It is worth noting that in Fig.˜4, CrisisSpot (w/o Att) refers to No Attention and CrisisSpot (with Attn) refers to the proposed method with IDEA attention mechanism. In summary, our model component analysis provides a nuanced understanding of CrisisSpot’s architecture, highlighting the critical role of each module in enhancing performance. Through systematic experimentation and insightful observations, we underscore the efficacy of our proposed approach in crisis content classification.
5.5.4.5 Analysis of Language, Vision and Graph Methods
In this section, we conduct a comprehensive analysis of language, vision, and graph models, focusing on their effectiveness in addressing crisis content classification tasks. Our selection of methods includes prominent representatives from each domain, namely BERT, ResNet, and GraphSAGE, each chosen for their distinct strengths and widespread adoption in the respective fields of natural language processing, computer vision, and graph representation learning, respectively.
BERT has demonstrated remarkable capability in capturing semantic nuances and contextual dependencies within textual data. Its superior performance across various natural language processing tasks, such as text classification, stems from its ability to leverage bidirectional context information effectively. Fig. 5 illustrates the accuracy, precision, recall, and F1-score achieved by various language and vision methods. Our analysis reveals that BERT outperforms both XLNet and RoBERTa across all the evaluation parameters. Similarly, ResNet is preferred for its depth and effectiveness in image classification, tackling vanishing gradient issues with skip connections and residual blocks. Its hierarchical feature capture across scales makes it a top choice for visual representation learning. ResNet when benchmarked against DenseNet and VGG-16 outperforms both the models with significant margins as clearly visible from Fig. 5. Furthermore, GraphSAGE was compared with GCN, in which GraphSAGE outperformed GCN across all evaluation metrics as depicted in Table 9. GraphSAGE is a pioneering method in graph representation learning that addresses the challenge of learning node embeddings in heterogeneous graph structures. By aggregating information from local neighborhoods through graph convolutions,
Figure 5: Analysis of Language and Vision Methods
GraphSAGE generates informative embeddings that encapsulate the structural and semantic properties of nodes in the graph. Its scalability, adaptability, and robustness make GraphSAGE a compelling choice for classification tasks. Our selection of BERT, ResNet, and GraphSAGE is driven by their proven effectiveness and widespread adoption in their respective domains.
Table 9: Analysis of Graph Methods
Method Accuracy Precision Recall F1
GCN 87.05 87.60 87.66 87.63
GraphSAGE 88.12 88.77 88.80 88.78
Numerical evidence, as depicted in the Fig. 5 and Table 9, further corroborates the superior performance of these baseline models compared to alternative methods across various classification benchmarks. CrisisSpot outperforms all of these methods across all evaluation metrics.
6 Discussion
This article presents CrisisSpot, a novel framework for classifying disaster-related content. CrisisSpot incorporates Social Context Features (SCFs) encompassing factors such as user post-informativeness, crisis post-informativeness, emotional context, and user engagement history, providing a multifaceted view of disaster-related data. Our proposed method introduces the Inverted Dual Embedded Attention (IDEA) mechanism, which adeptly directs attention to both aligned and misaligned information patterns across diverse modalities. Additionally, CrisisSpot employs a graph-based approach designed for the reconstruction of features which helps in enhancing the classification outcomes. Our proposed method leverages both cross-modal relationships and discriminative features to get a richer representation of the data. By leveraging these approaches, our work not only fills the gap in utilizing SCF but also enhances the model’s capability to effectively handle disaster-related information, thereby contributing significantly to the field of disaster content classification.
6.1 Advantages over Existing Works
Our proposed multimodal attentive framework offers several key advantages that elevate state-of-the-art methods for enhancing disaster content classification:
  • Integration of Social Context Features (SCF): A crucial aspect often neglected in prior works is the integration of SCF with the textual content. Our framework addresses this gap by introducing Social Context Features (SCF). SCF incorporates user-centric and text-centric information through metrics such as User Informative Score (UIS), Crisis Informative Score (CIS), and User Engagement metrics. UIS assesses the content informativeness based on their past posting history, while CIS quantifies the semantic relevance of the text content to the specific disaster event. These features empower the model to prioritize informative content posted by the users and content that is highly relevant to the crisis at hand.
  • Attention Mechanism: The proposed attention mechanism, Inverted Dual Embedded Attention (IDEA), focuses on both aligned and misaligned information patterns, potentially improving the model’s ability to handle complex multimodal data compared to conventional attention mechanisms limited to aligned information.
  • Graph-based Representation Learning: Many existing methods struggle to capture the intricate relationships between disaster-related content. Our framework addresses this limitation by integrating a graph-based approach. CrisisSpot utilizes cosine similarity and updates nodes using information from the entire dataset, leading to potentially more accurate feature refinement compared to existing methods limited by intra-cluster information or suboptimal distance metrics. CrisisSpot excels at learning local graph topology, crucial for capturing complex relationships, unlike knowledge graph-based methods that may face scalability challenges with large datasets.
  • Improved Classification Performance through Multimodal Learning: Most of the existing disaster content classification methods often rely solely on textual data. Our framework leverages a multimodal attentive approach, incorporating information from various modalities such as text and images. This fusion of modalities leads to a richer representation of disaster content, ultimately resulting in superior classification performance.
6.2 Limitations and Future Work
CrisisSpot demonstrates strong performance in disaster-related content classification. However, ongoing development can enhance its capabilities and broaden its applicability. The limitations of CrisisSpot are as follows:
  • Monolingual Focus: CrisisSpot currently focuses on monolingual data, achieving high accuracy within specific languages. This limits its effectiveness in multilingual disaster scenarios. Future work should explore techniques for incorporating multi-lingual capabilities. This could involve leveraging pre-trained multilingual transformer models or employing language identification modules to handle code-switching and mixed-language content.
  • Social Context Feature (SCF) Dependence: The inclusion of SCFs significantly enhances CrisisSpot’s classification performance. However, situations where SCF data might be limited necessitate exploring alternative feature extraction techniques.
  • Data Size and Augmentation: The size and quality of training data significantly impact performance. Exploring data augmentation techniques, such as back-translation, synonym substitution, or paraphrasing, could be a valuable avenue for further research. Data augmentation can potentially improve the model’s robustness by exposing it to a wider range of variations in disaster-related content.
6.3 Computational Efficiency and Scalability
CrisisSpot achieves significant advancements in disaster-related content classification while maintaining efficiency and scalability. This section details these aspects of our proposed approach.
Model Complexity and Efficiency: One key strategy to ensure computational efficiency is CrisisSpot’s utilization of three non-trainable encoders for feature extraction. These pre-trained encoders eliminate the need to train a significantly larger model from scratch, resulting in a total of 10.275 million trainable parameters. This parameter count falls within the range of typical deep learning architectures, mitigating potential computational burdens during training. CrisisSpot incorporates Social Context Features (SCF) that require minimal configuration, with only 21 neurons needed for SCF training. The proposed method demonstrates a real-time inference speed of 8.46 microseconds per sample.
Modular Architecture and Time Complexity: CrisisSpot leverages a modular architecture for efficient computation. The time complexity of each module is given as follows: Feature Extraction and Interaction: This module utilizes some operations such as scaled dot-product attention and a few matrix multiplications, contributing an overall time complexity of $O(mn^{2})$, where $m$ represents the maximum sequence length, and $n$ signifies the feature dimensionality of both textual and visual modalities. Multimodal Graph Learning Module: In this module, the complexity hinges on neighbor sampling and layer stacking, resulting in time complexity of $O(r^{k}n^{\prime}d^{2})$, where $r$ represents the fixed number of neighbors sampled per node, $k$ signifies the number of layers, $n^{\prime}$ represents the number of nodes, and $d$ represents the feature dimensions of both textual and visual modalities, which are fixed. Social Context Feature Module: This module operates in O(ij)O(i*j), where $i$ represents the input dimension and $j$ signifies the output dimension. However, since both $i$ and $j$ are small, the impact of this time complexity on the overall computation is minimal. The Multimodal Fusion Network comprises three networks with a combined eleven dense layers having a time complexity of $O(i^{\prime}*j^{\prime})$ for each sample where $i^{\prime}$ and $j^{\prime}$ denote the fixed input and output dimensions, respectively. This time complexity is minimal compared to other modules. The overall time complexity can be expressed as the sum of individual time complexities $O(mn^{2}+r^{k}n^{\prime}d^{2})$.
Scalability: It is also crucial to acknowledge that CrisisSpot, similar to other deep learning methods, employs a mini-batch training technique. This technique allows the model to process data in smaller batches during training. This approach offers two key advantages for scalability:
  • Efficient Handling of Varying Data Sizes: Mini-batch training enables CrisisSpot to efficiently process data across different scales, from smaller datasets to large-scale. Regardless of the volume of data, the model can be trained effectively while maintaining computational efficiency.
  • Flexibility for Distributed Training: Mini-batch training readily adapts to distributed computing environments. By distributing the training load across multiple machines or GPUs, CrisisSpot can be scaled to handle even larger datasets efficiently. CrisisSpot’s design is inherently scalable. The underlying transformer architecture can be readily parallelized on GPUs or TPUs for handling large datasets.
CrisisSpot prioritizes computational efficiency and scalability through its use of non-trainable encoders, real-time inference speed, and the adoption of mini-batch training. These design choices position CrisisSpot as a solution for disaster response, where accurate information is critical.
@@ 7 Conclusion
[…]